# What Should Embeddings Embed? Autoregressive Models Represent Latent Generating Distributions

### v1

@@ 4.3.2 Probing experiments
Implementational details
We use rectangles with seven unit blocks on each side, resulting in a size-784 hypothesis space, i.e., 784 different possible rectangles from which the sequences are drawn. The unit lengths are either 1 (Figure 5(a)), or 0.4 and 0.16 alternating (Figure 5(b)).
Probing results
We probe for the distribution $p(h|x_{1:n})$, which is a length-784 simplex vector, using the last token embedding of each sequence. Results are shown in Table 2. In general, the true hypothesis can be found with high accuracy, even though the number of classes is high.
Embedding visualization
We visualize the transformer embeddings to show what they represent in terms of this discrete set of hypotheses (Figure 6). We use Principal Components Analysis (PCA) to reduce the 128-dimensional embeddings to 2D, and color the points by properties of each rectangle. Embeddings are clustered into different regions based on the mean of the center of the true generating rectangle (Figure 6(a)), where, for instance, the top right yellow cluster corresponds to the upper right-most rectangle, and the middle teal cluster corresponds roughly to the many rectangles whose center have a mean of 3.5. Additionally, the embeddings encode the position along individual axes (Figure 6(b)), as well as distance between corners of the generating rectangle (Figure 6(c)).
(a) Colored by mean of the generating rectangle.
(b) Colored by mean along the y-axis of the generating rectangle.
(c) Colored by spread along the y-axis of the generating rectangle.
Figure 6: Two-dimensional representation of embeddings of all validation datapoints in the discrete hypothesis space dataset (the setup is unequal width and sample size $=50$). Subfigures show the same embeddings, colored by different properties of the true rectangle used to generate the data.
Table 2: Probing results on discrete hypothesis spaces. In general, the probe achieves strong performance in recovering a 784-length vector, and performance increases as task difficulty decreases. Average and standard deviation across 10 random seeds are reported.
Equal Width Unequal Width
20 $87.3\pm 1.2\%$ $0.173\pm 0.016$ $66\pm 1.7\%$ $0.29\pm 0.012$
50 $99.5\pm 0.2\%$ $0.008\pm 0.004$ $88.5\pm 0.9\%$ $0.159\pm 0.011$
@@ 5 Discussion
Limitations and Future Directions
Several limitations of our analysis motivate future work.
Extending to more complex Bayesian probabilistic models. We focused on three cases that consist of a one-step generation process and a simple likelihood $p(x|z)$ that can be approximated by a linear layer. Our analysis can be extended to hierarchical models or likelihoods parameterized by, for instance, neural networks, while simultaneously exploring more choices of transformer embeddings.
Interpreting neural networks with Bayesian probabilistic models. A similar Bayesian approach could be used to interpret the representations learned by other autoregressive neural networks. A direction for future work is extending our analyses to deep learning models for other modalities.
Task difficulty. As seen in the hypothesis space experiments, the ability to decode predictive sufficient statistics decreases as task difficulty increases. Developing a better understanding of the relationship between task difficulty and transformer embeddings is an important direction for future work.
Conclusion
We have developed a general framework for analyzing what the embeddings of an autoregressive language model should represent. Our analyses suggest that such embeddings should represent latent structures such that the next token $x_{n+1}$ is independent from previous tokens $x_{1:n}$ when conditioned on that structure, a property possessed by predictive sufficient statistics. We confirmed this hypothesis with probing experiments on three cases where predictive sufficient statistics can be identified. We hope that our findings contribute to bridging the gap between Bayesian probabilistic models and deep neural networks.
Acknowledgements
This work was supported by ONR grant number N00014-23-1-2510.
@@ Appendix A Appendix
A.1 Definition of additional exchangeable conjugate models
@@
A.1.1 Beta-Bernoulli model
@@
A.1.2 Gamma-Exponential model
@@
A.2 Implementational details
All training processes are run on a T4 or A100 Nvidia GPU.

### v2

@@ 4.3.1 Probing experiments
Experiment setup
The dataset is bags-of-words generated by LDA. We set the vocabulary size $V=10^{3}$, number of topics $K=5$, and generated $N=10^{4}$ documents that are each 100 words long.
Models
We trained four models: an autoregressive transformer decoder (AT), Bert, LDA, and an end-to-end word embedder (WE). Bert uses a masked objective instead of the autoregressive one, and is implemented by a small version called Bert-tiny (Turc et al., 2019). LDA is used to establish an upper-bound for model performance. We also include a model intended to provide an upper bound for embedding performance: a word embedder which is a matrix that maps from the vocabulary space to the AT / Bert embedding dimension and is end-to-end trained with the probe that predicts topics.
For each model, the hidden sizes and final-layer embedding sizes are 128. AT has 4 decoder layers to match the size of Bert-tiny. More hyperparameters are detailed in Appendix A.3.
Metrics
We use accuracy, L2 loss, and total variation loss to measure both classification performance and recovery of the topic mixture distribution. Accuracy is defined as how often the top topic predicted by the classifier’s mixture agrees with the top topic from ground truth. The remaining loss measures apply to the whole topic mixture.
Figure 6: Distribution of synthetic data topics predicted by LDA (row 1) and AT classifier (row 2) for different validation datapoints. Predictions are stick plots, and ground truth is bar chart in the background. They both exhibit learning of topic mixtures by trying to match the distribution, in addition to top-1 agreement.
Results
Figure 6 shows five examples of ground truth, transformer-predicted, and LDA-predicted topic mixtures. Table 3 shows that all models demonstrate success at recovering latent topics on at least the easiest setting (i.e., $\alpha=0.5$), being able to return both top-topic accuracy and the topic distribution spread. Between AT and Bert, the probe on AT is able to infer latent topic structures in more difficult tasks (i.e., $\alpha=0.8,1$) whereas the probe on Bert shows deteriorating performance. LDA outperforms both AT and Bert as expected because it is specified exactly to learn a dataset generated by the other manually initialized LDA. However, the strong WE performance suggests that the probe can predict topics mainly from stand-alone words. This raises the question whether AT and Bert are learning an underlying statistical model, or are simply uniquely embedding each word and making the probe mainly responsible for topic recovery.
4.3.2 Controlling for probe performance
In this section, we conduct control experiments that suggest that language models learn an underlying statistical model, making them—rather than the topic probes taking word embeddings —mainly responsible for successful topic recovery. If AT or Bert performs well just because it gives each word a unique embedding from which a trained probe suffices to recover the topic mixture, then a probe on top of the language model should additionally predict topic mixtures from a different underlying topic model than the one that generated the language model’s training data.
To assess this, we generate five datasets using five distinct topic models under the setting of $\alpha=0.5$. By definition, these topic models initialized five different sets of topics, or distributions over words $\{\beta_{1},...,\beta_{K}\}_{j},j\in\{1,2,3,4,5\}$. The transformer trained on set $j$ should encode the mixture of topics $\theta$ that corresponds to topics $\{\beta_{1},...,\beta_{K}\}_{j}$.
One AT and one Bert is trained on each dataset. On each model, five probes are used to predict topics from each of the five datasets. Results are shown in Figure 5. AT shows a strong distinction between predicting its own dataset versus datasets from other topic models, whereas this distinction is present but weaker for Bert. These results are evidence that AT and Bert indeed encode topic information because, if the topic information were instead constructed by the probe, the probe would work equally well on mismatched datasets as on matched ones.
4.4 From HMM-LDA to Natural Corpora
In this section, we start with a synthetic dataset with a more realistic assumption than topic models — that topics are an exchangeable component in a partially exchangeable sequence. We generate a dataset based on both LDA and a hidden Markov model (HMM). Then, we move to evaluate the extent to which LLMs recover topic mixture two natural corpora, 20Newsgroups (20NG) and WikiText-103.
Figure 7: Left: Illustration of two sequences (upper: $p\approx 0.5$; lower: $p\approx 0.25$). Each box is a word, coming from either the LDA-generated semantic component (blue), or one of the three categorical distributions that indicate three syntactic classes (grey). Right: Probing accuracy across semantic proportion $p$ on HMM-LDA generated datasets. Lower $\gamma$ corresponds to more sparse transition matrices. The language model successfully encodes semantics if it occupies a considerable enough proportion.
Table 4: 20NG topic prediction performance based on different LLMs. Trained LLMs substantially outperform the null GPT-2 model, supporting the hypothesis that the training process encourages LLMs to implicitly develop topic models. Autoregressive models (the first five models) statistically significantly outperform non-autoregressive models (the next two).
$K=20$ $K=100$
Model Parameters Accuracy $\uparrow$ L2 loss $\downarrow$ Tot. var. loss $\downarrow$ Accuracy $\uparrow$ L2 loss $\downarrow$ Tot. var. loss $\downarrow$
Gpt-2 124M $61.4\%\pm 1.5\%$ $0.106\pm 0.002$ $0.211\pm 0.001$ $42.3\%\pm 2.4\%$ $0.097\pm 0$ $0.192\pm 0.001$
Gpt-2-medium 355M $62.6\%\pm 1.7\%$ $0.104\pm 0.002$ $0.209\pm 0.002$ $42.9\%\pm 2.4\%$ $0.096\pm 0$ $0.19\pm 0.001$
Gpt-2-large 774M $62.4\%\pm 1.8\%$ $0.102\pm 0.002$ $0.208\pm 0.002$ $43.1\%\pm 2.3\%$ $0.095\pm 0.001$ $0.189\pm 0$
Llama 2 7B $62.6\%\pm 1.7\%$ $0.101\pm 0.002$ $0.206\pm 0.002$ $43.3\%\pm 2.4\%$ $0.095\pm 0.001$ $0.189\pm 0.001$
Llama 2-chat 7B $62.9\%\pm 1.7\%$ $0.102\pm 0.002$ $0.207\pm 0.002$ $43.2\%\pm 2.5\%$ $0.095\pm 0.001$ $0.189\pm 0$
Bert 110M $56.3\%\pm 1.5\%$ $0.113\pm 0.003$ $0.222\pm 0.003$ $38.6\%\pm 2.5\%$ $0.1\pm 0.001$ $0.191\pm 0.001$
Bert-large 336M $55.2\%\pm 1.2\%$ $0.116\pm 0.002$ $0.226\pm 0.003$ $38.9\%\pm 2.9\%$ $0.1\pm 0.001$ $0.191\pm 0.001$
Null Gpt-2 124M $27.3\%\pm 1\%$ $0.209\pm 0.003$ $0.322\pm 0.005$ $13.8\%\pm 1.7\%$ $0.145\pm 0.001$ $0.248\pm 0.003$
4.4.1 Topic distribution can be recovered when LDA is a sequence sub-component
Most natural texts are not exchangeable. For example, consider this sentence: “The sediment found in the quartz includes silicon.” This sentence is neither fully exchangeable nor generated purely by a latent state model. However, if we only consider the words that contribute to the sentence’s topic (geology), we are left with sediment, quartz, silicon. These words can be plausibly generated by an exchangeable topic model as a sub-component: the sentence “The silicon found in the sediment includes quartz” also has the topic of geology even though word order changes. While we do not exhaustively list possible factors behind language that can be embedded by LLMs, we study to what extent can topics be encoded by model embeddings under settings where topical words occupy a varied proportion with respect to the whole sequence.
Model
Similar to the LDA experiments, we generate synthetic documents with an HMM-LDA model (Griffiths et al., 2004) (Figure 7). HMM-LDA combines semantics and syntax, and posits that each word comes from a latent class and that class transitions are governed by an HMM. Using the HMM-LDA model allows one to manipulate the degree to which topic exists as a sub-component of the sequence.
When $p$ is 1, all the words are generated from the LDA model. As $p$ decreases, more words are generated from the syntactic classes. We can then examine how well topic distributions are recovered from the embeddings of models that are trained on data generated from HMM-LDA models that vary $p$.
Results
Results in Figure 7 suggest that without exchangeability, the autoregressive transformer still learns the same semantic latent variables provided there is a large enough proportion of LDA-generated component (i.e., $p\geq 0.4$). This suggests that the theory of finding explainable quantities in autoregressive models can potentially be extended to real texts where interpretable probabilistic models, such as the topic model, exist as a sub-component of language.
4.4.2 Recovering latent topic distributions in natural corpora
This section turns to analyzing LLMs pretrained on natural language. Although an exhaustive list of predictive sufficient statistic is difficult to identify, topical words do form a proportion of factors underlying the text. This allows us to hypothesize that topic mixtures can be decoded from the transformer residual streams.
Datasets
We use 20Newsgroups (20NG) and WikiText-103 (Merity et al., 2016). 20NG is a collection of 18,000 posts written in a style similar to informal emails, divided into 20 subjects. Contrasting with the informal language style in 20NG, WikiText-103 consists of over 100 million tokens sourced from the set of verified articles on Wikipedia that are classified as Good and Featured.
Setup
We train LDA models across three random seeds on each dataset, and use pretrained large language models as our LLM. Specifically, the LLMs are Gpt-2, Gpt-2-medium, and Gpt-2-large (Radford et al., 2019); Llama 2 and Llama 2-chat (Touvron et al., 2023); and Bert, and Bert-large (Devlin et al., 2019). The probes target LDA-learned topic mixtures with $K\in\{20,100\}$. For the models, we additionally include a randomly initialized Gpt-2, called Null Gpt-2, as a control that can differentiate the contribution of the LLM from the contribution of the probe. Considering that the first token may lead to a useful embedding that stands for the $\langle\text{CLS}\rangle$ token in Bert series, we searched across the first, last, and average embeddings.
Topic prediction results
Performance is shown in Table 4 for 20NG and in Table 7 in the Appendix for Wikitext-103. To predict mixtures of twenty topics, random guessing would yield a $5\%$ accuracy for $K=20$ or $1\%$ accuracy for $K=100$. The best-performing LLM on each dataset demonstrates success at encoding topic distributions by achieving $88.5\%/74.2\%$ accuracy on WikiText-103, and $62.9\%/43.2\%$ on 20NG. The probes here take averaged token embeddings on the last layer in each LLM. While having direct information over whole sequences is advantageous compared to over an arbitrary token, we show that using only the last token preserves strong performance, with $73.7\%/58.9\%$ accuracy on WikiText-103, and $52.2\%/34.7\%$ on 20NG (Table 14 and 13 in Appendix). These results suggest that topic mixtures are directly encoded in embeddings, rather than resulting from high-quality word representations.
Topics encoded by inner layers
We also explored the possibility of decoding topic mixtures from other LLM layers, as opposed to only the last layer (Table 8 in Appendix A.5). In the Llama-2 series, the decodability of topic mixtures progressively increases as we move from the word embedder to the intermediate layers. Although a word embedder can potentially achieve high accuracy as demonstrated by the synthetic dataset, we observe that Llama-2 relies on its inner layers to encode topic mixtures. This phenomenon can be caused by Llama-2’s stronger ability to model the next word, which is correlated with the ability to capture latent structures, which we discuss in more detail in the next paragraph.
Probe performance and LLM perplexity
To show that the the posterior predictive on topics corresponds to autoregressive prediction, we analyze 100 probes trained on 100 different token positions, along with their corresponding LLM perplexity (Figure 5(c)). Each position is defined based on the corresponding percentile of the total document (each document has a different length). We expect that as perplexity on a token increases, probe performance based on the embedding taken from that token would decrease. This hypothesis is supported by the linear trend in Figure 5(c). A possible confound in this result is that perplexity tends to decrease as the position in a document increases, so the results could be driven by position rather than perplexity; however, we ran a linear mixed-effects model and found that perplexity continued to have a statistically significant effect controlling for position (see Appendix A.4).
@@ 5 Discussion
In this paper, we have shown that autoregressive modeling is connected to Bayes optimal agents, where the sufficient statistic is encoded by the model in two major cases: 1) exchangeable distributions that emcompass a wide range of generative models including the topic model; and 2) latent state models that include the HMM. We have also seen that when these latent variables form a significant subcomponent of the data, they can still be encoded, leading to an explanatory approach on why topic mixtures are embedded in transformer residual streams.
There are several potential implications and future directions of our findings.
Interpretability.
Understanding the inner workings of LLMs is important for AI safety and trustworthiness. However, the sheer size of LLMs makes it challenging to analyze. Our results suggest that interpretable models of document structure - such as topic models - provide useful guidance about what mechanistic interpretability should look for. This in turn means that one goal for interpretability work should be enhancing models like LDA because such advances will in turn sharpen our ability to interpret LLM representations.
Reverse direction - training LLMs to capture inductive biases.
We have shown how autoregressive models capture latent structures when predicting the next token. Another direction to explore is whether it is possible to explicitly incorporate inductive biases during training to more reliably and efficiently capture these latent structures, as have been done in modeling language (McCoy and Griffiths, 2025) and vision (Carballo-Castro et al., 2024).
Extensions to masked language models.
It has been shown that the autoregressive objective can be broken down into the Bayesian predictive distribution in several cases. It is possible to extend this breakdown to masked language models as well (Appendix A.1). Whereas the autoregressive model learns a different posterior distribution $p(\theta|x_{1:n})$ for each $n$ as it progresses through tokens in a sequence, the masked model learns that same $p(\theta|x_{U})$, where $U$ denotes the set of unmasked indices, resulting in a less expressive objective. On synthetic topic model bags-of-words, the masked model underperforms, but it is worth exploring whether its performance is consistently worse than that of the autoregressive model in a wider range of cases.
Applications.
Our results help inform construction and use of embeddings in practice, and can be applied to generative models that satisfy the conditions discussed by the paper. The analyses can be naturally extended to other latent variables that do not depend heavily on word order, such as the author type of the document (Andreas, 2022) or the author’s sentiment (Radford et al., 2017).
As another example, our results can also inform time series modelling. A practitioner might have a time series dataset with underlying factors that are informed by human experts. If the practitioner wants to use deep autoregressive models to construct an embedding that contains a certain kind of information from the input, they may need to ensure that this information is a predictive sufficient statistic for the task.
6 Conclusion
We have developed a general framework for analyzing what the embeddings of an autoregressive language model should represent. Our analyses suggest that such embeddings should represent latent structures such that the next token $x_{n+1}$ is independent from previous tokens $x_{1:n}$ when conditioned on that structure, a property possessed by predictive sufficient statistics. We confirmed this hypothesis with probing experiments on cases where predictive sufficient statistics can be identified. We hope that our findings contribute to bridging the gap between Bayesian probabilistic models and deep neural networks.
Acknowledgments
This work was supported by ONR grant number N00014-23-1-2510. TRS is supported by an NDSEG Fellowship. RTM was supported by the NSF SPRF program under Grant No. 2204152.
@@ Appendix A Appendix
A.1 Masked language models
Masked language models (MLMs) are another class of LLMs that have been successful at modeling text, and we explore whether the objective of MLMs is also equivalent to implicit Bayesian inference under the exchangeability assumption. We find that it is, but it differs from autoregressive language models in a way that results in less expressivity.
MLMs, notably including the Bert series (Devlin et al., 2019), randomly mask certain tokens in the input sequence, and the model’s goal is to predict the original words at the masked indices only by considering their surrounding context. Unlike autoregressive models, the MLM does not model a coherent joint distribution of the data (Yamakoshi et al., 2022; Young and You, 2022). However, the log objective can be extended as follows,
$\mathcal{L}_{MLM}(x_{1:N})$ $=\sum_{n\in M}\log p(x_{n}|x_{i,i\in U})$
$=\sum_{n\in M}\log\int p(x_{n}|\theta)p(\theta|x_{i,i\in U})d\theta.$ (4)
where $M$ denotes the set of masked indices, and $U$ denotes the set of unmasked indices. The proof is given below, and is a simple extension of the derivation for the autoregressive version. The difference between the MLM objective and the autoregressive objective is that in the summation, the prediction of each token $x_{n}$ uses the same posterior over the latent variable $p(\theta|x_{i,i\in U})$. In other words, each token $x_{n}$ is predicted independently from the latent variable $\theta$. As a result, MLM forms a less expressive Bayesian inference objective than autoregressive models. Therefore, we aim to empirically evaluate both the ability of MLM to recover latent variables, and whether its performance differs from that of autoregressive models.
We now prove the equivalence between the MLM objective and Bayesian inference in Equation 4. The statement is that, given an exchangeable process $x_{1},x_{2},...,x_{n}$,
$\mathcal{L}_{MLM}(x_{1:N})$ $:=\sum_{n\in M}\log p(x_{n}|x_{i,i\in U})$ (5)
$=\sum_{n\in M}\log\int p(x_{n}|\theta)p(\theta|x_{i,i\in U})d\theta,$ (6)
where $M$ is the set of masked indices and $U$ is the set of unmasked indices, and by construction $M\cap U=\emptyset$.
Proof.
We first prove the autoregressive version of the equivalence, which Korshunova et al. (2018) proposes for each individual term in the summation but briefly mentions why it is equivalent,
$\log p(x_{1:N})$ $=\log p(x_{1})+\sum_{n=1}^{N-1}\log p(x_{n+1}|x_{1:n})$
$=\log p(x_{1})+\sum_{n=1}^{N-1}\log\int_{\theta}p(x_{n+1}|\theta)p(\theta|x_{1:n})d\theta.$ (7)
To do so, we shall prove the equivalence in each term in the summation, that is, the statement that
$p(x_{n+1}|x_{1:n})=\int_{\theta}p(x_{n+1}|\theta)p(\theta|x_{1:n})d\theta$ (8)
for each $n>1$. First, de Finetti’s theorem states that, under the same exchangeability condition,
$p(x_{1:n+1})=\int_{\theta}p(\theta)\prod_{i=1}^{n+1}p(x_{i}|\theta)d\theta,$ (9)
and also that each $x_{i}$ is conditionally independent given $\theta$ for all $i$. To show the equivalence, we first divide each side of Equation 9 by $p(x_{1:n})$, assuming that $p(x_{1:n})>0$. The left hand side becomes,
$\frac{p(x_{1:n+1})}{p(x_{1:n})}=p(x_{n+1}|x_{1:n}).$ (10)
The right hand side becomes,
$\int_{\theta}p(\theta)\frac{\prod_{i=1}^{n+1}p(x_{i}|\theta)}{p(x_{1:n})}d\theta$ $=\int_{\theta}p(\theta)\frac{p(x_{n+1}|\theta)p(x_{1:n}|\theta)}{p(x_{1:n})}d\theta$ (11)
$=\int_{\theta}p(x_{n+1}|\theta)\frac{p(\theta)p(x_{1:n}|\theta)}{p(x_{1:n})}d\theta$ (12)
$=\int_{\theta}p(x_{n+1}|\theta)p(\theta|x_{1:n})d\theta.$ (13)
Line 11 uses conditional independence to combine the product on $i=1$ through $n$, and line 13 uses Bayes rule. Therefore, because of Equation 9, we prove the statement of Equation 8.
The proof for MLM in Equation 6 can be shown by a simple extension. We shall use $x_{U}$ as short hand for $x_{i,i\in U}$. Using de Finetti’s theorem,
$p(x_{\{n\}\cup U})=\int_{\theta}p(\theta)\prod_{i\in\{n\}\cup U}p(x_{i}|\theta)d\theta,$ (14)
Dividing each side of the above equation by $p(x_{U})$, the left hand side becomes,
$\frac{p(x_{\{n\}\cup U})}{p(x_{U})}=p(x_{n}|x_{U}).$ (15)
The right hand side becomes,
$\int_{\theta}p(\theta)\frac{\prod_{i\in\{n\}\cup U}p(x_{i}|\theta)}{p(x_{U})}d\theta$ $=\int_{\theta}p(\theta)\frac{p(x_{n}|\theta)p(x_{U}|\theta)}{p(x_{U})}d\theta$ (16)
$=\int_{\theta}p(x_{n}|\theta)\frac{p(\theta)p(x_{U}|\theta)}{p(x_{U})}d\theta$ (17)
$=\int_{\theta}p(x_{n}|\theta)p(\theta|x_{U})d\theta.$ (18)
Same as in the proof for the autoregressive version, line 16 uses conditional independence to combine the product on $U$, and line 18 uses Bayes rule. Therefore, we have that $p(x_{n}|x_{i,i\in U})=\int p(x_{n}|\theta)p(\theta|x_{i,i\in U})d\theta$. Thus, we have shown the equivalence for each term in the summation of the original statement Equation 6.
A.2 Definition of additional exchangeable conjugate models
@@
A.2.1 Topic model
The generative model for LDA is
For each topic $k$ in (1,,K),(1,...,K), 1. Draw topic $\beta_{k}\sim\text{Dirichlet}_{V}(\eta)$. For each document $i$, 1. Draw topic mixture $\theta_{i}\sim\text{Dirichlet}(\alpha).$ 2. For each word $j$ in document $i$, (a) Draw topic assignment $t_{ij}\sim\text{Categorical}(\theta_{i})$, (b) Draw word $x_{ij}\sim\text{Categorical}(\beta_{t_{ij}})$,
where $V$ is the vocabulary size, and $\alpha$ and $\eta$ are pre-initialized hyperparameters.
A.2.2 Beta-Bernoulli model
@@
A.2.3 Gamma-Exponential model
@@
A.2.4 HMM-LDA
There are $C$ classes $\{c_{1},...,c_{C}\}$ which follow an HMM transition matrix. One class is the semantic class $c_{1}$, and a word coming from $c_{1}$ is generated by an LDA model. Words of other classes are ‘syntactic’ and are directly generated by a distribution over words with a Dirichlet prior, one distribution for each $\{c_{2},...,c_{C}\}$. The probability of generating from $c_{1}$, the LDA class, is determined by a parameter $p$ which specifies the transition probability into that class from all other classes. The remaining transition probabilities are generated from a Dirichlet($\gamma$) prior; lower $\gamma$ corresponds to more sparse transition matrices.
A.3 Implementational details
All computations for synthetic datasets are run on single Tesla T4 GPUs, and those for natural corpora are run on single A100 GPUs.